OpenStudy (unofficialllyy):

At the movie theatre, child admission is $6.30 and adult admission is $9.90. On Sunday, 168 tickets were sold for a total sales of $1393.20. How many adult tickets were sold that day?

OpenStudy (unofficialllyy):

@Kurayami95

OpenStudy (kurayami95):

6.30a+9.90=1393.20 a+b=168

OpenStudy (kurayami95):

6.30a+9.90b=1393.20

OpenStudy (kurayami95):

use that equation intead i forgot to add b ._. sorry

OpenStudy (unofficialllyy):

thats okay, thanks!!